OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. A(5, 5), m = 3

OpenStudy (anonymous):

y = 3x + c To find c, the poin A(5, 5) is on the line, SO 5 = 3*5 + c ==> c = 5 -15 = -10 The required line is : y = 3x - 10

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.